OpenStudy (anonymous):

of what is 3/7 of 21

OpenStudy (anonymous):

9

OpenStudy (anonymous):

can you tell me how you got 9 please

OpenStudy (anonymous):

3/7 of 21: The "mathematical version" of "of" is "x" or "times" So: 3/7*21 =21*3/7 =63/7 =9

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

what is your real name